# VideoCAD: A Large-Scale Video Dataset for Learning UI Interactions and 3D Reasoning from CAD Software

### v1

@@ Appendix A Onshape
We use Onshape [5], a cloud-native CAD platform, as the environment for our task of learning CAD UI for creating and editing 3D models of mechanical parts and assemblies. Note that unlike traditional desktop-based CAD software, Onshape runs entirely in the browser and supports real-time collaboration, version control (similar to Git), and parametric design. We chose OnShape as it is free, platform agnostic (a majority of CAD software run on Windows), and highly accessible via the browser. In addition, OnShape contains most of the CAD operations used in standard commercial CAD software. We choose to use Google Chrome as the browser of choice as it is available on all main operating systems.
Figure 8: Overview of the Onshape user interface. The feature bar at the top includes general modeling operations as well as sketch-specific tools such as lines, rectangles, and circles. The left panel displays the feature tree with the default geometry (Top, Front, Right planes), while the central view shows the coordinate system and sketching planes. The right panel demonstrates options for the Extrude feature, used after defining sketches.
A.1 DeepCAD Representation
The DeepCAD sequence [4] represents modeling operations as a sequence of the form:
$[t_{i},x,y,\alpha,f,r,\theta,\phi,\gamma,p_{x},p_{y},p_{z},s,e_{1},e_{2},u,b]$
Each segment corresponds to a command: - ti: command type (0: line, 1: arc, 2: circle, 4: loop separator, 5: extrusion) - x, y: points - $\alpha,f$: arc angle and direction - r: radius (for circle) - $\theta,\phi,\gamma$: plane angles - $p_{x},p_{y},p_{z}$: sketch origin offset - s: sketch scale - e1, e2: extrusion extents - u, b: extrusion operation type and symmetry
A.2 Normalization
All values from the DeepCAD sequences are in $[0,255]$ and are normalized using:
$\mathcal{N}(p)=\frac{p-128}{128}$
Used for coordinates, angles, plane offsets, and extrusion parameters.
A.3 Plane Basis and Extrusion Plane Parameters
The sketch plane is defined by normalized angles $(\theta,\phi,\gamma)$ and a 3D offset $(p_{x},p_{y},p_{z})$. Using:
$(\theta,\phi,\gamma)=\pi\cdot\mathcal{N}([\theta,\phi,\gamma]),\quad\mathbf{o}=\mathcal{N}([p_{x},p_{y},p_{z}])$
we compute an orthonormal basis $(\mathbf{n},\mathbf{x},\mathbf{y})$, where:
$\mathbf{n}=\text{polar-to-cartesian}(\theta,\phi,\gamma),\quad\mathbf{y}=\mathbf{n}\times\mathbf{x}$
The sketch plane is mapped to an integer ID:
[…]
The final extrusion plane offset is:
[…]
A.4 Point Transformation to Pixel Space
Every geometric point in a sketch (e.g., endpoints, centers, midpoints) is transformed into 2D pixel coordinates through the following steps:
  1. 1.
    Normalization:
    $\mathbf{p}_{\text{norm}}=\mathcal{N}([x,y])=\frac{[x,y]-128}{128}$
  2. 2.
    Rotation and Offset:
    $\mathbf{p}_{\text{rot}}=R\cdot\mathbf{p}_{\text{norm}}\cdot s+\mathbf{o}$
    where $R=[\mathbf{x},\mathbf{y}]^{T}$ is the rotation matrix derived from the sketch plane basis, $s$ is the global scale, and $\mathbf{o}$ is the origin offset.
  3. 3.
    Projection: We project the rotated 3D point to 2D by dropping the axis corresponding to the sketch plane’s normal direction.
    Define a binary vector $\texttt{mask}\in\{0,1\}^{3}$ that keeps the two in-plane components:
    $\texttt{mask}[i]=\begin{cases}0&\text{if }i=\texttt{plane\_id}\\ 1&\text{otherwise}\end{cases}$
    The 2D projected point becomes:
    $\mathbf{p}_{\text{proj}}=\mathbf{p}_{\text{rot}}[\texttt{mask}]$
  4. 4.
    Pixel Alignment:
    $\mathbf{p}_{\text{pixel}}=0.5\cdot\mathbf{p}_{\text{proj}}+\mathbf{C}$
    where $\mathbf{C}$ is the pixel-space origin of the Onshape canvas.
This transformation is consistently applied to all sketch entities including:
  • Line endpoints
  • Circle centers
  • Arc midpoints, center, start and end points
A.5 Primitive-Specific Parameter Computation
Each loop in the sketch is composed of geometric primitives: lines, circles, and arcs, each defined by a set of transformed parameters. Below, we define each primitive mathematically using pixel-space coordinates.
A.5.1 Lines
A line segment is defined by its start and end points. Let $\mathbf{p}_{\text{start}},\mathbf{p}_{\text{end}}\in\mathbb{R}^{2}$ be the 2D pixel coordinates after projection:
$\texttt{Line}=\left\{\mathbf{p}_{\text{start}},\mathbf{p}_{\text{end}}\right\}$
Each point is obtained by:
$\mathbf{p}=0.5\cdot\left(R\cdot(\mathcal{N}([x,y])\cdot s)+\mathbf{o}\right)[\texttt{mask}]+\mathbf{C}$
A.5.2 Circles
A circle is defined by its center and radius. Let $\mathbf{c}\in\mathbb{R}^{2}$ be the projected center and $r$ the scaled radius:
$r=\frac{r_{\text{seq}}}{128}\cdot s\cdot 0.5$
$\texttt{Circle}=\left\{\mathbf{c},r\right\},\quad\text{with }\mathbf{c}=0.5\cdot\left(R\cdot(\mathcal{N}([x,y])\cdot s)+\mathbf{o}\right)[\texttt{mask}]+\mathbf{C}$
A.5.3 Arcs
An arc is defined by its start point $\mathbf{p}_{\text{start}}$, end point $\mathbf{p}_{\text{end}}$, center $\mathbf{c}_{\text{arc}}$, midpoint $\mathbf{p}_{\text{mid}}$, and radius $r$.
Angle and Direction.
$\alpha=180\cdot\frac{\alpha_{\text{seq}}}{128},\quad f\in\{0,1\}$
Chord Geometry.
$\mathbf{v}=\mathbf{p}_{\text{end}}-\mathbf{p}_{\text{start}},\quad L=\|\mathbf{v}\|,\quad\mathbf{p}_{\text{chord}}=\frac{\mathbf{p}_{\text{start}}+\mathbf{p}_{\text{end}}}{2}$
Radius and Offset.
$r=\frac{L}{2\sin(\alpha/2\cdot\pi/180)},\quad h=\sqrt{r^{2}-(L/2)^{2}}$
Center and Midpoint.
$\mathbf{v}_{\perp}=\frac{[-v_{y},v_{x}]}{\|\mathbf{v}\|},\quad\mathbf{c}_{\text{arc}}=\begin{cases}\mathbf{p}_{\text{chord}}+h\cdot\mathbf{v}_{\perp},&\text{if }f=1\\ \mathbf{p}_{\text{chord}}-h\cdot\mathbf{v}_{\perp},&\text{if }f=0\end{cases}$
The angular midpoint is computed as:
$\theta_{\text{start}}=\text{atan2}(p_{\text{start},y}-c_{y},p_{\text{start},x}-c_{x})$
$\theta_{\text{mid}}=\theta_{\text{start}}\pm\frac{\alpha}{2}\cdot\frac{\pi}{180}$
$\mathbf{p}_{\text{mid}}=\mathbf{c}_{\text{arc}}+r\cdot\begin{bmatrix}\cos(\theta_{\text{mid}})\\ \sin(\theta_{\text{mid}})\end{bmatrix}$
Representation.
$\texttt{Arc}=\left\{\mathbf{p}_{\text{start}},\ \mathbf{p}_{\text{end}},\ \mathbf{c}_{\text{arc}},\ \mathbf{p}_{\text{mid}},\ r\right\}$
All points are transformed to pixel coordinates using the transformation pipeline from Section LABEL:appendix:cad_generation.
A.6 Extrusion Parameters
Extrusion parameters are computed as:
[…]
$e_{1}=0.5\cdot\mathcal{N}(e_{1}),\quad e_{2}=0.5\cdot\mathcal{N}(e_{2})$
A.7 Final Representation
Each extrusion is encoded as:
[…]
Where profile is a list of loops, and each loop is a list of geometric primitives: lines, arcs, and circles with parameters as computed above.
Appendix B Detailed Onshape UI Action Procedure
This appendix provides a thorough description of our automated CAD modeling procedure within Onshape, driven by our hybrid UI interaction framework based on rule-based commands. The procedure systematically translates structured CAD command sequences (defined in Appendix LABEL:appendix:cad_generation) into executable UI interactions within Onshape.
B.1 Plane Creation Procedure
For sketches requiring custom-defined planes (offset from default planes), we follow these UI steps:
  1. 1.
    Activate the plane creation tool by clicking the plane icon.
  2. 2.
    Select one of the default reference planes (Top, Front, Right).
  3. 3.
    Navigate to the offset text box via successive presses of the Tab key.
  4. 4.
    Enter the desired offset value numerically.
  5. 5.
    Click the offset direction arrow to define plane offset orientation.
  6. 6.
    Finalize plane creation by pressing Enter.
B.2 Sketch Creation and Loop Building Procedure
Each sketch comprises loops defined by geometric primitives: lines, arcs, and circles. These loops are drawn via the following UI actions:
General Sketch Setup:
  1. 1.
    Start a new sketch by pressing Shift+S.
  2. 2.
    Select the appropriate sketch plane (custom or default).
Drawing Geometric Primitives:
  • Line: Press L, then click to specify start and end points.
  • Circle: Press C, click to set the center point, move cursor outward to specify radius, then click to finalize.
  • Arc (3-point): Press A, sequentially click to set start, end, and midpoint.
Constraint Management:
To ensure geometric accuracy, constraints are toggled dynamically:
  • Press Shift key down to deactivate constraints.
  • Press Shift key up to activate constraints (primarily when connecting loop endpoints).
Command Completion:
  • Exit current geometric command (line, arc, circle) using Esc.
B.3 Visibility and Navigation
Visibility and navigation are managed to maintain clarity and prevent interference from previously drawn elements:
  • Toggle visibility of planes: Shift+P.
  • Toggle visibility of sketches: Shift+H.
  • Hide/show previously built parts: Y/Shift+Y.
  • Navigate between default planes (Top, Front, Right) using key sequences: Shift (down)+(Up/Right/Down/Left)Shift (up).
B.4 Extrusion Execution Procedure
Extrusions convert sketch loops into 3D features through these UI steps:
  1. 1.
    Select sketch region(s) intended for extrusion.
  2. 2.
    Activate extrusion using Shift+E.
  3. 3.
    Choose extrusion type (New for adding material, Remove for material removal).
  4. 4.
    Navigate (using Tab) to set the extrusion depth numerically.
  5. 5.
    Navigate further (using Tab) to the symmetric option checkbox; press Space to toggle symmetry on/off.
  6. 6.
    If performing material removal (Remove), navigate to and activate the “Merge with all” checkbox.
  7. 7.
    Finalize extrusion by pressing Enter.
B.5 End-of-Sequence Token
We use the hotkey combination Shift+7 to set the camera view to an isometric perspective. This action signifies the completion of the CAD modeling sequence and serves as our end-of-sequence token.
Appendix C Ablation on Vision Models for CAD Image Similarity
Table 6: Comparison of CLIP and DINOv2 for CAD image similarity over 400 samples
Metric CLIP DINOv2
Mean of correct scores 0.8576 0.7923
Median of correct scores 0.8587 0.8022
Std Dev of correct scores 0.0250 0.0488
Mean of incorrect scores 0.7931 0.6002
Median of incorrect scores 0.7947 0.6014
Std Dev of incorrect scores 0.0167 0.0440
Average rank of correct match 18.24 2.21
Median rank of correct match 6.00 1.00
Std Dev of rank 31.69 3.08
We conduct an ablation comparing DINOv2 and CLIP as image encoders for CAD shape comparison. For each of 400 query samples, we compute cosine similarity between features extracted from a rendered UI isometric image and a set of 400 ground-truth isometric CAD images (including the correct one). We define the correct score as the similarity between the UI image and its associated CAD image, and the incorrect scores as the similarities with all other CAD images in the set. The rank of the correct match is computed by ranking all 400 similarity scores in descending order.
Table 6 reports the results. Although CLIP yields slightly higher correct match scores on average, its incorrect scores are also high—indicating limited discriminative power for geometric retrieval. In contrast, DINOv2 achieves significantly lower similarity scores for incorrect matches (mean of 0.6002 vs. 0.7931 for CLIP), resulting in more robust separation and better retrieval accuracy. This is reflected in the rank-based metrics: DINOv2 achieves an average correct match rank of 2.21 and a median of 1.00, far outperforming CLIP (18.24 and 6.00, respectively), with far less variance (std dev of 3.08 vs. 31.69).
This improvement can be attributed to the architectural and training differences between the models: CLIP is optimized for semantic alignment between text and image—learning “what is this”—whereas DINOv2 is trained in a self-supervised manner to model fine-grained visual structure, capturing both local and global geometry. This makes DINOv2 naturally better suited for comparing CAD-like images, where shape and spatial configuration are more critical than semantic labels.
Threshold Selection. We also use these similarity scores to define a threshold for dataset quality control. Specifically, we set the threshold at 0.7, which corresponds to the average of the median correct score (0.8022) and median incorrect score (0.6014). This ensures a conservative yet robust filter for identifying high-quality image-CAD pairs.
Appendix D Dataset Benchmarking
The description of the metrics used to compare dataset complexity is explained in the following: # Samples: number of samples in the dataset, Time Horizon: number of UI interactions needed to complete the task, Requires 3D Reasoning: we look through every app used in each dataset. We determine whether the app requires the agent to reason about 3D coordinates to manipulate UI-state. Contains Precise Element: many benchmarks ground agents by leveraging the Document Object Model (DOM), enabling agents to select UI elements without needing to specify xy coordinates. However, some elements such as canvas elements cannot be interacted via the DOM. We define a precise element as an element that requires an agent to manipulate via xy coordinates, which require a higher level of reasoning as agents must rely on spatial and visual reasoning skills to click on the correct buttons. To check whether a benchmark requires a precise element, we look through every app used in each dataset and determine whether the agent is required to interact with a precise element such as a canvas, Average UI elements: To show that OnShape’s complexity as a GUI, we measure the number of elements as described in Mind2Web. We only report the average number of elements for datasets that include the HTML tree in the dataset.
Appendix E Training Details and Ablation Studies
Training Details
@@ Appendix E Training Details and Ablation Studies
All models were trained using 4 NVIDIA H100 GPUs. VPT and VideoCAD required 4 days of training, while Pix2Act and Pearce et al. converged in 1 day. Quantitative results are summarized in Table E.
We train VideoCADFormer on the VideoCAD dataset with a 90/5/5 train/val/test split. At each step, the model autoregressively predicts the next action given a window of $k=10$ past UI frames and actions. Outputs are treated as classification tasks and trained using cross-entropy loss with inverse-frequency class weighting. For pointer and typed parameters ($x$, $y$, $v$), we apply soft targets over a ±$\pm 2$ bin window to account for geometric tolerance.
Table 7: Ablation results showing the impact of input types, image modalities, and architecture choices on command accuracy, parameter accuracy, and perfect action prediction across task difficulties
@@
Finally, we explore VideoCADFormer’s ability to generalize to different image types such as realistic images and sketches (Figure 9). We first generated images from a DeepCAD sequence using a geometry kernel. From this kernel-generated image, we constructed photo-realistic images by passing the kernel-generated CAD images into ControlNet [35] . The sketches are generated by applying canny line detection to the kernel-generated CAD images and then applying a Gaussian blur. Among the three, the model performs best on sketches, followed by kernel-generated images, and performs worst on photorealistic images, suggesting increased abstraction aids generalization.
@@
Appendix F Metrics
@@
Appendix G Generated samples
@@
Appendix H Failure Analysis
@@
Appendix I Benchmarking LLMs’ 3D Reasoning Capabilities
@@
Table 8: Questions from our CAD-VQA dataset
@@
Collectively, these questions exercise geometric recognition, viewpoint awareness, forward simulation, temporal planning, and numeric inference — topics that are rarely interrogated simultaneously by existing VQA benchmarks. We generated 200 samples for each question for a total of 2,200 samples, though this question generation process can be easily scaled to the entire dataset. We provide the full benchmarking results on the questions above in Table 5.2.

### v2

@@ Appendix A CAD Software
We use Onshape [onshape], a cloud-native CAD platform, as the environment for our task of learning CAD UI for creating and editing 3D models of mechanical parts and assemblies. Note that unlike traditional desktop-based CAD software, Onshape runs entirely in the browser and supports real-time collaboration, version control (similar to Git), and parametric design. We chose Onshape as it is free, platform agnostic (a majority of CAD software run on Windows), and highly accessible via the browser. In addition, Onshape contains most of the CAD operations used in standard commercial CAD software. We choose to use Google Chrome as the browser of choice as it is available on all main operating systems.
Figure 7: Onshape user interface
Appendix B Model Architecture and Ablation Studies
Input Representation.
Visual Encoding.
$z_{t}^{\text{image}}=\phi_{\text{proj}}([v^{I};v^{o}_{t}])\in\mathbb{R}^{h}.$
Action and Timestep Embeddings.
Transformer Decoder.
We use two attention masks: a causal mask $\mathbf{M}^{\text{causal}}\in\mathbb{R}^{T\times T}$ is an upper-triangular matrix with -$\infty$ in positions where future tokens should be masked and 0 elsewhere
Action Prediction.
$\hat{c}_{t}$ $=\text{softmax}(W_{c}H_{t}+b_{c}),\qquad\hat{c}_{t}\in\mathbb{R}^{5},\qquad\qquad\hat{p}_{t}=\text{softmax}(W_{p}H_{t}+b_{p}),\qquad\hat{p}_{t}\in\mathbb{R}^{6\times 1000}.$
Command-dependent masks $M_{\hat{c}_{t}}\in\{0,1\}^{6}$ are applied to suppress unused parameter outputs. Invalid values are set to 1-1.
@@ Appendix B Model Architecture and Ablation Studies
All models were trained using 4 NVIDIA H100 GPUs. VPT and VideoCAD required 4 days of training, while Pix2Act and Pearce et al. converged in 1 day. Quantitative results are summarized in Table B.
We train VideoCADFormer on the VideoCAD dataset with a 90/5/5 train/val/test split. At each step, the model autoregressively predicts the next action given a window of $k=10$ past UI frames and actions. Outputs are treated as classification tasks and trained using cross-entropy loss with inverse-frequency class weighting. For pointer and typed parameters ($x$, $y$, $v$), we apply soft targets over a $\pm 2$ bin window to account for geometric tolerance.
Table 6: Ablation results showing the impact of input types, image modalities, and architecture choices on command accuracy, parameter accuracy, and perfect action prediction across task difficulties.
@@
Finally, we explore VideoCADFormer’s ability to generalize to different image types such as realistic images and sketches (Figure 8). We first generated images from a DeepCAD sequence using a geometry kernel. From this kernel-generated image, we constructed photo-realistic images by passing the kernel-generated CAD images into ControlNet [zhang2023addingconditionalcontroltexttoimage] . The sketches are generated by applying canny line detection to the kernel-generated CAD images and then applying a Gaussian blur. Among the three, the model performs best on sketches, followed by kernel-generated images, and performs worst on photorealistic images, suggesting increased abstraction aids generalization.
@@
Appendix C Evaluation Metrics for CAD Generation
@@
Appendix D Uncertainty Quantification and Structural Fidelity of VideoCADFormer
Measured Uncertainty Validates Significance of Results. To quantify uncertainty, we estimate the standard error of the success rate using: σ= p(1 - p)n,   n = 200
For the Medium-length category:
VPT: $\qquad p=0.364\Rightarrow\sigma\approx 3.4\%$
VideoCADFormer: $\qquad p=0.379\Rightarrow\sigma\approx 3.4\%$
VideoCADFormer Excels in Structural Fidelity and Robustness. The modest difference between VPT and VideoCADFormer in medium-length sequences falls within statistical uncertainty, even after increasing the sample size. However, our model consistently outperforms across short and long sequences, achieving lower Chamfer Distance and generating fewer invalid CAD models across all sequence lengths.
We hypothesize that VPT tends to produce sequences that initially resemble the ground truth, performing adequately on medium-complexity cases. As sequence complexity increases (Long category), however, VPT fails to preserve structure and parametric correctness, leading to sharp drops in geometric fidelity. In contrast, VideoCADFormer maintains spatial and parametric consistency over longer horizons, demonstrating stronger generalization and robustness.
Appendix E Image-Conditioned Generated CAD models using VideoCADFormer
@@
Appendix F Failure Analysis of VideoCADFormer
@@
Appendix G Conducting a Failure Analysis on LLM Agents on CAD Design Tasks
Our current failure analysis of LLM agents is qualitative, as all of the tested agents were unable to make meaningful progress during evaluation. We have enclosed several thought trace samples to better illustrate the observed failure modes. As these models improve, we anticipate that more quantitative benchmarking will become meaningful.
We identify two common failure modes: (1) a lack of domain understanding of Onshape, and (2) grounding or perception issues.
Addressing Domain Understanding Challenges
For issues pertaining to limited Onshape understanding, we commonly observe traces such as the following:
Action: It looks like my previous attempt to click the "Restore" button failed
because it might not have been correctly identified. Given that the current
context is focused on the "Sketch 1" with the message that it has been canceled,
the next logical step is to see if I can find the "Restore" button again or use
another strategy to re-engage with the sketch. I’ll check for available buttons
or features related to the sketch in the current interface. Since there’s no
immediate visible button to "Restore," I might try searching for features related
to the sketch or navigating through the interface. Next, I will scroll to see if
more options or buttons become visible.
scroll(0, 100)
In this case, the model attempted to restore a sketch by identifying alternative UI elements. However, it failed to recognize that the correct action was simply Ctrl+Z, instead issuing a scroll command on the sketch plane of the UI.
Addressing Grounding Challenges
For issues related to grounding, we observe traces such as:
Action: Given the current context and the recurring issues with previous actions,
its important to strategize the next step carefully:
1. Challenges Encountered: There have been errors when attempting to click
specific features, indicating either hidden elements or overlapping actions.
2. Current Focus: I’m in the Onshape part studio, where the workspace is largely
empty, with no parts created and the "Origin" feature hidden.
3. Goal: I need to find a way to effectively manipulate the sketch or geometry.
Strategy:
To move forward, it seems prudent to start by resetting my focus within the
workspace. Ensuring that my interaction is centered correctly might open up
further possibilities for manipulating features.
Next Action:
I will attempt to click on the "Sketch" section to see if that allows me to
access or reveal any relevant tools or options moving forward.
click("34")
Here, the model attempts to select the sketch section by referencing an internal element ID. However, in this case, "34" corresponds to a navigation bar element, not the intended sketch section, highlighting a fundamental grounding failure between perceived and actual interface elements.
Appendix H Benchmarking LLMs’ 3D Reasoning Capabilities
@@
Table 7: Questions from our CAD-VQA dataset.
@@
[…]